OpenStudy (latinrose):

Using the formula r=d/t, where d is the distance in miles, r is the rate, and t is the time in hours, at which rate must you travel to to cover 337.5 miles in 4.5 hours?

OpenStudy (agent0smith):

miles is distance. So that gives you d. hours is time. So that gives you t. Plug them in.

OpenStudy (latinrose):

I literally have dyscalculia so I have no idea

OpenStudy (agent0smith):

you know r=d/t you have d = 337.5 t = 4.5 so \[\large r = \frac{ 337.5 }{ 4.5 }\]

OpenStudy (latinrose):

The question I'm answering though is asking for an answer in mph though? How do I translate \[r=\frac{ 337.5 }{ 4.5 }\] into mph?

OpenStudy (agent0smith):

By working it out. You are dividing miles by hours. Thus, the units are miles/hour.

OpenStudy (agent0smith):

\[\large r=\frac{ 337.5 \text{ miles}}{ 4.5 \text{ hours}} = ? \text{ miles/hour }\]